### v1

# Intrinsic Diophantine approximation: a solution to Mahler’s problem

Abstract.
In the 80’s Mahler suggested to study the Diophantine approximation by rationals lying in the middle-third Cantor set $K_{1/3}$. In this article we address this question by proving that, for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+},$
$\dim_{H}\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }\frac{p}{q}\in\mathbb{Q}\cap K_{1/3}\right\}=\frac{\dim_{H}K_{1/3}}{\max\left\{1,\delta_{\psi}\right\}},$
where $\delta_{\psi}$ is the shrinking rate of $\psi.$ This result will be derived out a more general theorem regarding homogeneous self-similar rational IFS’s. In addition, we establish several results regarding the intrinsic Diophantine approximation associated with inhomogeneous rational IFS’s.
1. Introduction
The study of rational approximation on self-similar fractals was first suggested by Mahler [21]. He asked how well numbers in the middle-third Cantor set can be approximated by rationals, and also by rationals lying in the middle-third Cantor set itself. Both of these questions have attracted much interest, and many results have recently been established.
Among these results, Benard, He and Zhang [3] showed that if $\mu$ is a self-similar measure on $\mathbb{R}$ (Definition 2.3), then for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$ one has (where “f.i.m.” stands for “for infinitely many”)
$\mu\Big(\Big\{x\in\mathbb{R}:\ |x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\Big\}\Big)=1\ \Leftrightarrow\ \sum_{q\geq 1}q\psi(q)=+\infty,$
$\mu\Big(\Big\{x\in\mathbb{R}:\ |x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\Big\}\Big)=0\ \Leftrightarrow\ \sum_{q\geq 1}q\psi(q)<+\infty.$
Concerning the dimension of sets of points approximable at rate $\delta\geq 2$ (that is, when $\psi(q)=q^{-\delta}$, the rate $2$ being the threshold provided by Dirichlet’s theorem) by rational numbers in the middle-third Cantor set $K_{1/3}$, Bugeaud and Durand [7] conjectured that
$\dim_{H}\Big\{x\in K_{1/3}:\ |x-\tfrac{p}{q}|\leq q^{-\delta}\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{1/3}\Big\}$
$=\max\Big\{\frac{\dim_{H}K_{1/3}}{\delta},\ \dim_{H}K_{1/3}+\frac{2}{\delta}-1\Big\}.$
Here, the first term corresponds to the intrinsic contribution (obtained by considering the approximation by rationals lying in the middle-third Cantor set), while the second corresponds to the extrinsic one. In this direction, Chow, Yu and Varjú [9] established that, in the case of the middle-seventh Cantor set, there exists $\delta_{0}>2$ such that the above formula holds for any $2\leq\delta\leq\delta_{0}$. This result was later generalized by Chen (using techniques from [3]) to any self-similar set satisfying the open set condition [8].
Regarding intrinsic approximation, namely approximation by rationals lying inside a given self-similar set, it is worth mentioning that it is not generic for a self-similar set to contain rational numbers. Indeed, given a homogeneous self-similar set $K\subset\mathbb{R}$, it is easily checked that for every $y\in\mathbb{R}$, $y+K$ is also the attractor of a homogeneous self-similar IFS. Moreover, one can show that
[…]
The interested reader may refer to the appendix for a proof of this fact.
A natural class of self-similar sets known to contain rational numbers consists of those defined by rational parameters. The first results in this setting were obtained by Beresnevich and Velani [4], and refined by Levesley, Salp and Velani [20], who provided a complete description of the metric theory of $\psi$-approximable points in base 3 in $K_{1/3}$. The problem was also studied, among others, by Fishman and Simmons [18], Tan, Wang and Wu [24], and Baker [1]. Writing $h(r)$ the height of $r\in\mathbb{Q},$ that is, $h(r)=q$ if $r=\frac{p}{q}$ with $p\wedge q=1$, in [18] and [24], the authors proposed the following conjecture (expressed differently), extending that of Bugeaud and Durand: for any non-increasing mapping $\psi:\mathbb{N}\to\mathbb{R}_{+}$, writing $\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q},$ [^1] A careful reader will notice that the set investigated by Bugeaud and Durand does not involve the height $h(r)$. It as small exercise to show that, for any non increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$ $\left\{x\in K_{1/3}:\ |x-\frac{p}{q}|\leq\psi(q)\text{ f.i.m. }(p,q)\in\mathbb{Z}\times\mathbb{N}\text{ s.t. }\frac{p}{q}\in K_{T}\right\}=\left\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\text{ f.i.m. }r\in\mathbb{Q}\cap K_{T}\right\}\bigcup(\mathbb{Q}\cap K_{T})$. Thus, the choice of the definition of the approximating sets does not change the dimension results.
$\dim_{H}\Big\{x\in K_{1/3}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{1/3}\Big\}=\frac{\dim_{H}K_{1/3}}{\max\Big\{1,\ \delta_{\psi}\Big\}}.$
We confirm this conjecture by proving the following result.
Theorem 1.1.
Let $b\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$ be integers, and set
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\Big\}_{1\leq i\leq m}.$
Let $K_{T}$ be the attractor of $T$ (see Definition 2.3). If $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, writing
$\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q},$
we have
$\dim_{H}E_{\psi,T}:=\dim_{H}\Big\{x\in K_{T}:\ |x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}}.$
Notice that, in the above formula, $\frac{\dim_{H}K_{T}}{\max\left\{1,\delta_{\psi}\right\}}=\min\left\{\dim_{H}K_{T},\frac{\dim_{H}K_{T}}{\delta_{\psi}}\right\}.$ Moreover, it is a routine check to verify that
$\frac{\dim_{H}K_{T}}{\delta_{\psi}}=\inf\left\{s\geq 0:\ \sum_{n\geq 1}e^{n\dim_{H}K_{T}}\psi\Big(e^{n}\Big)^{s}<+\infty\right\}.$
In contrast, by Khintchine’s theorem, when $K_{T}$ has non empty interior, one has, for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+},$
$\dim_{H}E_{\psi,T}=\min\left\{1,s_{\psi}\right\},$
where
$s_{\psi}=\inf\left\{s\geq 0:\ \sum_{n\geq 1}n\psi(n)^{s}<+\infty\right\}.$
To illustrate this striking dichotomy, we state the following corollary regarding the intrinsic Jarnik sets.
Corollary 1.2.
Let $b\in\mathbb{N}$ and $p_{1},....,p_{m}\in\mathbb{Z}$ be integers and set
$T=\left\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\right\}_{1\leq i\leq m}.$
Let $K_{T}$ be the attractor of $T$, $\delta\geq 1$ and write $\psi_{\delta}:q\mapsto q^{-\delta}.$ Then
[…]
The second purpose of this article is to provide more general results for inhomogeneous rational IFSs. Given integers $m\geq 2$, $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, for $1\leq i\leq m$ define, for any $x\in\mathbb{R},$
(1) $f_{i}(x)=\frac{x}{q_{i}}+\frac{p_{i}}{q_{i}},$
let $T=\{f_{1},\dots,f_{m}\}$ and denote by $K_{T}$ its attractor. We recall the following fact.
Lemma 1.3 ([18, 1]).
Let $r\in\mathbb{Q}$. Then $r\in K_{T}$ if and only if there exist $\underline{i}=(i_{1},\dots,i_{n})\in\bigcup_{p\geq 1}\{1,\dots,m\}^{p}$ and $0\leq k\leq n-1$ such that
$r=\pi\Big((i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty})\Big),$
where $\pi$ is the canonical projection associated with $T$.
A straightforward computation shows that there exists $p\in\mathbb{Z}$ such that
$\pi\Big((i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty})\Big)=\frac{p}{\prod_{1\leq j\leq k}q_{i_{j}}\,\big(\prod_{k+1\leq j\leq n}q_{i_{j}}-1\big)}.$
One of the main difficulties in intrinsic Diophantine approximation is that there is, in general, no obvious way to reduce this fraction. Consequently, in [18, 1] the authors introduced a natural intrinsic height for $r\in\mathbb{Q}\cap K_{T}$, which is in general larger than the usual height, by setting
$h_{\mathrm{int}}(r)=\min\Big\{\prod_{1\leq j\leq k}q_{i_{j}}\,\big(\prod_{k+1\leq j\leq n}q_{i_{j}}-1\big)\Big\},$
where the minimum is taken over all representations of $r$ of the above form. With this definition, the natural set to consider is
$E_{\psi,T,\text{int}}=\Big\{x\in K_{T}:\ |x-r|\leq\psi(h_{\mathrm{int}}(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\}.$
In [24], Tan, Wang and Wu provided a complete metric description of $E_{\psi,T,\text{int}}$ for the middle-third Cantor set, and in [1], Baker gave conditions under which these sets have full measure with respect to the natural self-similar measure on $K_{T}$. In the present article, we provide a complete description of the Hausdorff dimension of the sets $E_{\psi,T,\text{int}}$.
Theorem 1.4.
Let $T$ be the IFS defined by (1), $K_{T}$ its attractor, and let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be non-increasing. Then
$\dim_{H}E_{\psi,T,\text{int}}=\frac{\dim_{H}K_{T}}{\max\Big\{1,\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}\Big\}}.$
It is worth mentioning that we do not require any separation condition on $T$. In particular, it is not known in general whether $\mathcal{H}^{\dim_{H}K_{T}}(K_{T})>0$ for IFSs of the form (1). Moreover, this result follows from a more general theorem concerning approximation by ultimately periodic sequences on self-similar fractals (see Theorem 4.1).
Finally, as for any $r\in\mathbb{Q}\cap K_{T},$ $h_{\text{int}}(r)\geq h(r)$, the following corollary holds.
Corollary 1.5.
Let $T$ be a self-similar IFS defined by maps $f_{i}$ as in (1). Then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$,
$\dim_{H}\Big\{x\in K_{T}:\ |x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{T}\Big\}\geq\frac{\dim_{H}K_{T}}{\max\left\{1,\liminf_{n\to+\infty}\frac{-\log\psi(n)}{\log n}\right\}}.$
2. Notations, recalls and main statements
Let us start with some notation
Let $d$ $\in\mathbb{N}$. For $x\in\mathbb{R}^{d}$, $r>0$, $B(x,r)$ stands for the closed ball of ($\mathbb{R}^{d}$,||||||\ \ $|_{\infty}$) of center $x$ and radius $r$. Given a ball $B$, $|B|$ stands for the diameter of $B$. For $t\geq 0$, $\delta\in\mathbb{R}$ and $B=B(x,r)$, tBtB stands for $B(x,tr)$, i.e. the ball with same center as $B$ and radius multiplied by $t$, and the $\delta$-contracted ball $B^{\delta}$ is defined by $B^{\delta}=B(x,r^{\delta})$.
[…]
Given a topological space $X$, the Borel $\sigma$-algebra of $X$ is denoted $\mathcal{B}(X)$ and the space of probability measures on $\mathcal{B}(X)$ is denoted by $\mathcal{M}(X).$
The $d$-dimensional Lebesgue measure on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ is denoted by $\mathcal{L}^{d}$.
For $\mu\in\mathcal{M}(\mathbb{R}^{d})$, $\mathrm{supp}(\mu)=\left\{x\in\mathbb{R}^{d}:\ \forall r>0,\ \mu(B(x,r))>0\right\}$ is the topological support of $\mu$.
Given $E\subset\mathbb{R}^{d}$, $\dim_{H}(E)$ and $\dim_{P}(E)$ denote respectively the Hausdorff and the packing dimension of $E$.
2.1. Recalls on geometric measure theory and fractal geometry
We now recall some definitions.
Definition 2.1.
Let $\zeta:\mathbb{R}^{+}\to\mathbb{R}^{+}$ be increasing in a neighborhood of 0. The Hausdorff outer measure at scale $t\in(0,+\infty]$ associated with $\zeta$ of a set $E$ is defined by
(2) $\mathcal{H}^{\zeta}_{t}(E)=\inf\left\{\sum_{n\in\mathbb{N}}\zeta(|B_{n}|):\ |B_{n}|\leq t,\ B_{n}\text{ closed balls and }E\subset\bigcup_{n\in\mathbb{N}}B_{n}\right\}.$
The Hausdorff measure associated with $\zeta$ of a set $E$ is defined by
$\mathcal{H}^{\zeta}(E)=\lim_{t\to 0^{+}}\mathcal{H}^{\zeta}_{t}(E).$
For $t\in(0,+\infty]$, $s\geq 0$ and $\zeta(x)=x^{s}$, we write $\mathcal{H}^{\zeta}_{t}(E)=\mathcal{H}^{s}_{t}(E)$ and $\mathcal{H}^{\zeta}(E)=\mathcal{H}^{s}(E)$, which are respectively called the $s$-dimensional Hausdorff outer measure at scale $t$ and the $s$-dimensional Hausdorff measure. Thus,
(3) $\mathcal{H}^{s}_{t}(E)=\inf\left\{\sum_{n\in\mathbb{N}}|B_{n}|^{s}:\ |B_{n}|\leq t,\ B_{n}\text{ closed balls and }E\subset\bigcup_{n\in\mathbb{N}}B_{n}\right\}.$
The quantity $\mathcal{H}^{s}_{\infty}(E)$ is called the $s$-dimensional Hausdorff content of $E$.
Definition 2.2.
Let $\mu\in\mathcal{M}(\mathbb{R}^{d})$. For $x\in\mathrm{supp}(\mu)$, the lower and upper local dimensions of $\mu$ at $x$ are defined by
$\underline{\dim}_{\rm loc}(\mu,x)=\liminf_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r},\qquad\overline{\dim}_{\rm loc}(\mu,x)=\limsup_{r\to 0^{+}}\frac{\log\mu(B(x,r))}{\log r}.$
The lower Hausdorff dimension and upper packing dimension of $\mu$ are respectively defined by
(4) $\dim_{H}(\mu)=\mathrm{ess\,inf}_{\mu}(\underline{\dim}_{\rm loc}(\mu,x)),\qquad\overline{\dim}_{P}(\mu)=\mathrm{ess\,sup}_{\mu}(\overline{\dim}_{\rm loc}(\mu,x)).$
It is known (see [14] for details) that
$\dim_{H}(\mu)=\inf\{\dim_{H}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\ \mu(E)>0\},$
$\overline{\dim}_{P}(\mu)=\inf\{\dim_{P}(E):\,E\in\mathcal{B}(\mathbb{R}^{d}),\ \mu(E)=1\}.$
When $\dim_{H}(\mu)=\overline{\dim}_{P}(\mu)$, this common value is denoted by $\dim_{H}(\mu)$ and $\mu$ is said to be exact dimensional.
2.2. Self-similar iterated function systems
2.2.1. Basic definitions and facts
A map $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a contracting similarity if there exist $\lambda\in\mathbb{R}\setminus\left\{0\right\}$ with $|\lambda|<1$, $O\in\mathcal{O}_{d}(\mathbb{R})$ and $a\in\mathbb{R}^{d}$ such that
$f(x)=\lambda Ox+a.$
The number $c=|\lambda|$ is called the contraction ratio of $f$.
Definition 2.3.
A self-similar IFS is a family $S=\{f_{i}\}_{1\leq i\leq m}$ of $m\geq 2$ contracting similarities of $\mathbb{R}^{d}$. Let $(p_{i})_{1\leq i\leq m}$ be a probability vector.
The associated self-similar measure $\mu$ is the unique probability measure satisfying
(5) $\mu=\sum_{i=1}^{m}p_{i}\,\mu\circ f_{i}^{-1}.$
Its support is the attractor $K$, i.e. the unique non-empty compact set such that
$K=\bigcup_{i=1}^{m}f_{i}(K).$
The existence and uniqueness of $K$ and $\mu$ are standard [19], and by a result of Feng–Hu [16], self-similar measures are exact dimensional. Moreover, it was established by Falconer in [13] that any self-similar set $K$ satisfies
$\dim_{H}K=\dim_{B}K,$
where $\dim_{B}$ denotes the box dimension.
Given a self-similar IFS $T=\{f_{1},\dots,f_{m}\}$, we write
[…]
2.2.2. Separation properties of rational IFSs
Let $T=\{f_{1},\dots,f_{m}\}$ be a self-similar IFS, let $K$ be its attractor, and let $0<c_{1}\leq\cdots\leq c_{m}<1$ be the contraction ratios of $f_{1},\dots,f_{m}$. For $r>0$, we define
$\Lambda_{r}=\Big\{\underline{i}=(i_{1},\dots,i_{n})\in\bigcup_{k\geq 1}\{1,\dots,m\}^{k}:\ c_{\underline{i}}\leq r<c_{(i_{1},\dots,i_{n-1})}\Big\}.$
We begin by recalling the definition of the Asymptotically Weak Separation Condition (AWSC). This property was introduced by Feng in [15] and applies to a very large class of IFSs (see [2]), as mentioned in the introduction.
Definition 2.4.
We say that $T$ satisfies the AWSC if
$\lim_{r\to 0^{+}}\frac{\log\Big(\sup_{x\in K}\#\big\{f_{\underline{i}}:\ \underline{i}\in\Lambda_{r}\text{ and }f_{\underline{i}}(K)\cap B(x,r)\neq\emptyset\big\}\Big)}{\log r}=0.$
The following result is proved in [2, Theorem 1.5].
Theorem 2.1 ([2]).
Let $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$. Then the IFS
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\Big\}_{1\leq i\leq m}$
satisfies the AWSC.
When the above IFS $T$ is homogeneous, i.e. when $q_{1}=\cdots=q_{m}$, it satisfies a stronger separation condition.
Definition 2.5.
Let $T=\{f_{1},\dots,f_{m}\}$ be a self-similar IFS. We say that $T$ satisfies the Weak Separation Condition (WSC) if
$\sup_{x\in K,\ r>0}\#\big\{f_{\underline{i}}:\ \underline{i}\in\Lambda_{r}\text{ and }f_{\underline{i}}(K)\cap B(x,r)\neq\emptyset\big\}<+\infty.$
For references regarding IFS’s satisfying the weak separation conditions, see for instance [17, 25]. Given $q\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, it is straightforward to check that
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{q}+\frac{p_{i}}{q}\Big\}_{1\leq i\leq m}$
satisfies the WSC. Moreover, Zerner established the following result.
Theorem 2.2 ([25]).
Let $T$ be a self-similar IFS satisfying the WSC and let $K_{T}$ be its attractor. Then
$0<\mathcal{H}^{\dim_{H}K_{T}}(K_{T})<+\infty.$
Although it is not stated explicitly, it is a direct consequence of Theorem 2.2 that if $T$ satisfies the WSC, then $K_{T}$ is Ahlfors regular. For the sake of completeness, we prove this fact in the appendix.
Proposition 2.3.
Let $T$ be a self-similar IFS satisfying the WSC. Then $K_{T}$ is $\dim_{H}K_{T}$-Ahlfors regular.
3. Proof of the upper bound in Theorem 1.1
Let $m,b\in\mathbb{N}$, let $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\Big\}.$
Denote by $K_{T}$ the attractor of $T$. In this section, we prove that if $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$ such that
$\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}\leq 1,$
writing
$E_{\psi,T}=\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K_{T}\Big\},$
one has
$\dim_{H}E_{\psi,T}\leq\frac{\dim_{H}K_{T}}{\delta_{\psi}}.$
Together with Theorem 1.4 (see Corollary 1.5), this proves Theorem 1.1.
Let $\widetilde{K}_{T}\subset\mathbb{T}^{1}$ be the projection of $K_{T}$ onto the one-dimensional torus. It is easily verified that $x\in\widetilde{K}_{T}$ implies $bx\in\widetilde{K}_{T}$. Moreover, $\widetilde{K}_{T}$ is compact; hence, if $\widetilde{K}_{T}\neq\mathbb{T}^{1}$, then $\mathbb{T}^{1}\setminus\widetilde{K}_{T}$ is a non-empty open set, while if $\widetilde{K}_{T}=\mathbb{T}^{1}$, then $K_{T}$ has non-empty interior.
We will prove the following result.
Theorem 3.1.
Let $b\in\mathbb{N}$ and let $S\subset\mathbb{T}^{1}$ be a $\times b$-invariant set. For every $n\in\mathbb{N}$, set
$Q_{b,n}=\Big\{b^{n-1}<q\leq b^{n}:\ \exists p\in\mathbb{N},\ \frac{p}{q}\in S,\ p\wedge q=1\Big\}.$
Assume that $\mathbb{T}^{1}\setminus S$ has non-empty interior. Then
$\lim_{n\to+\infty}\frac{\log\#Q_{b,n}}{n}=0.$
Before proving Theorem 3.1, we explain how it implies Theorem 1.1.
Let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be non-increasing. Observe that
$E_{\psi,T}$ $=\Big\{x\in K_{T}:|x-\tfrac{p}{q}|\leq\psi(q)\ \text{f.i.m. }\tfrac{p}{q}\in\mathbb{Q}\cap K_{T},\ p\wedge q=1\Big\}.$
Lemma 3.2.
Let $q\in\mathbb{N}$. Then there exists a constant $C>0$ such that
$\Big\{p\in\mathbb{N}:\ \frac{p}{q}\in K_{T}\Big\}\leq Cq^{\dim_{H}K_{T}}.$
Proof.
For $p_{1}\neq p_{2}$ with $\frac{p_{1}}{q},\frac{p_{2}}{q}\in K_{T}$, one has
$\Big|\frac{p_{1}}{q}-\frac{p_{2}}{q}\Big|\geq\frac{1}{q},$
so the balls $B(\frac{p_{1}}{q},\frac{1}{2q})$ and $B(\frac{p_{2}}{q},\frac{1}{2q})$ are pairwise disjoint and centered in $K_{T}$. Since, by Proposition 2.3, $K_{T}$ is Ahlfors regular, there exists $\mu\in\mathcal{M}(K_{T})$, a $\dim_{H}K_{T}$-Ahlfors regular probability measure supported on $K_{T}$. Then
$\sum_{p:\ \frac{p}{q}\in K_{T}}\mu\Big(B\Big(\frac{p}{q},\frac{1}{2q}\Big)\Big)\leq 1,$
which implies the desired bound, with $C$ depending only on $\mu$. ∎
We are now ready to establish the upper bound. For any $N\in\mathbb{N}$,
$E_{\psi,T}\subset\bigcup_{n\geq N}\ \bigcup_{\begin{subarray}{c}q\in Q_{b,n}\\ p:\ \frac{p}{q}\in K_{T}\end{subarray}}B\Big(\frac{p}{q},\psi(b^{n})\Big).$
Assume that Theorem 3.1 holds. Fix $\varepsilon>0$ and choose $N_{\varepsilon}$ large enough so that for every $k\geq N_{\varepsilon}$ and every $\widetilde{k}\geq\frac{\log N_{\varepsilon}}{\log b}$,
[…]
Set
$s_{\varepsilon}=\frac{\dim_{H}K_{T}+\varepsilon}{\delta_{\psi}}+\varepsilon.$
Then, by Lemma 3.2, there exists $\kappa>0$ such that
$\sum_{q\geq N_{\varepsilon}}\sum_{p:\ \frac{p}{q}\in K_{T}}\psi(q)^{s_{\varepsilon}}$ $\leq\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}+1}\kappa\,b^{n\dim_{H}K_{T}}\,b^{\varepsilon n}\,b^{-n(1-\varepsilon)\delta_{\psi}s_{\varepsilon}}$
$=\kappa\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n(\delta_{\psi}s_{\varepsilon}-\dim_{H}K_{T}-\varepsilon)}$
$\leq\kappa\sum_{n\geq\frac{\log N_{\varepsilon}}{\log b}}b^{-n\varepsilon\delta_{\psi}}<+\infty.$
Hence $\dim_{H}E_{\psi,T}\leq s_{\varepsilon}$, and letting $\varepsilon\to 0$ yields the claim. Before proving Theorem 3.1, we recall basic definitions and results related to discrepancy of sequences on $\mathbb{T}^{1}$.
3.1. Recall on equidistribution
We begin by recalling the definition of discrepancy.
Definition 3.1.
Let $U\subset\mathbb{T}^{1}$ be a finite set with $U=N<+\infty$. The discrepancy of $U$ is defined by
$D(U)=\sup_{0<a<b<1}\Big|\frac{\#\{u\in U:\ u\in[a,b]\}}{N}-(b-a)\Big|.$
Remark $3.$
From the definition, it is immediate that if $D(U)\leq\eta$ and $J\subset\mathbb{T}^{1}$ is an interval with $\mathcal{L}(J)>\eta$, then there exists $u\in U$ such that $u\in J$.
We also recall the Erdős–Turán inequality.
Theorem 3.4 ([12]).
Let $U$ be a finite sequence with #U=N\#U=N. Then, for any $A\in\mathbb{N}$,
$D(U)\leq C\Big(\frac{1}{A}+\sum_{1\leq j\leq A}\frac{1}{Nj}\Big|\sum_{u\in U}\exp(2\pi iju)\Big|\Big).$
3.2. Proof of Theorem 3.1
We fix $b$ and $S$ as in the theorem, and denote by $\ell>0$ the length of an interval $I\subset\mathbb{T}^{1}\setminus S$ By considering a larger constant $C$ in Theorem 3.4 if one must, we may assume that $\frac{C}{2\ell}\in\mathbb{N}$.
Given $q\in\mathbb{N}$, we write $\mathbb{Z}_{q}=\mathbb{Z}/q\mathbb{Z}$ and $\mathbb{Z}_{q}^{*}$ for its group of invertible elements. For $a\in\mathbb{Z}_{q}^{*}$, we denote by $o_{q}(a)$ the multiplicative order of $a$ modulo $q$.
We first recall the following result due to Bourgain.
Theorem 3.5 ([5]).
Let $q\in\mathbb{N}$ and let $H$ be a subgroup of $\mathbb{Z}_{q}^{*}$. For any $\varepsilon>0$ there exists $\eta_{\varepsilon}>0$ such that, provided that $q$ is large enough,
(6) $H\geq q^{\varepsilon}\quad\Longrightarrow\quad\max_{p\in\mathbb{Z}_{q}^{*}}\Big|\sum_{h\in H}\exp\Big(2\pi i\frac{hp}{q}\Big)\Big|\leq q^{-\eta_{\varepsilon}}\#H.$
We start with a lemma.
Lemma 3.6.
Let $\varepsilon>0$. Then for any $n\in\mathbb{N}$ and any $q\in Q_{b,n}$ with $q\wedge b=1$ and $q$ large enough, there exists $1\leq t\leq\frac{2C}{\ell}$ such that
$o_{q/(q\wedge t)}(b)<\Big(\frac{q}{q\wedge t}\Big)^{\varepsilon}.$
Proof.
Assume on the contrary that for every $1\leq t\leq\frac{2C}{\ell}$,
$o_{q/(q\wedge t)}(b)\geq\Big(\frac{q}{q\wedge t}\Big)^{\varepsilon}.$
Since $q\in Q_{b,n}$, there exists $p\in\mathbb{Z}_{q}^{*}$ such that $\frac{p}{q}\in S$. As $S$ is $\times b$-invariant, this implies that
$U=\Big\{u_{k}:=\frac{b^{k}p}{q}\Big\}_{0\leq k\leq o_{q}(b)-1}\subset S.$
Applying the Erdős–Turán inequality (Theorem 3.4) with $A=\frac{2C}{\ell}$, we obtain
$D(U)\leq\frac{\ell}{2}+\sum_{1\leq h\leq\frac{2C}{\ell}}\frac{1}{ho_{q}(b)}\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|.$
For $1\leq h\leq\frac{2C}{\ell}$, write $q^{\prime}=\frac{q}{h\wedge q}$ and $h^{\prime}=\frac{h}{h\wedge q}$. Then
$\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\Big|\frac{o_{q}(b)}{o_{q^{\prime}}(b)}\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|.$
Hence
$\frac{1}{o_{q}(b)}\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|.$
Since $o_{q^{\prime}}(b)\geq{q^{\prime}}^{\varepsilon}$, by Theorem 3.5,
$\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|\leq{q^{\prime}}^{-\eta_{\varepsilon}}.$
Therefore,
$D(U)\leq\frac{\ell}{2}+\frac{2C}{\ell}\,q^{-\eta_{\varepsilon}}.$
For $q$ large enough this implies $D(U)<\ell$, hence there exists $0\leq k\leq o_{q}(b)-1$ such that $u_{k}\in I\subset\mathbb{T}^{1}\setminus S$, yielding a contradiction. ∎
Proposition 3.7.
For all sufficiently large $n\in\mathbb{N}$ and every $\varepsilon>0$,
$N_{\varepsilon,n,b}=\#\Big\{b^{n-1}<q\leq b^{n},\ b\wedge q=1:\ \exists\,1\leq t\leq\frac{2C}{\ell}\text{ with }o_{q/(q\wedge t)}(b)\leq q^{\varepsilon}\Big\}\leq b^{2\varepsilon n}.$
This follows from the next lemma.
Lemma 3.8.
For all sufficiently large $n\in\mathbb{N}$ and every $\varepsilon>0$,
$\Big\{b^{n-1}<q\leq b^{n},\ b\wedge q=1:\ o_{q}(b)\leq q^{\varepsilon}\Big\}\leq b^{2\varepsilon n}.$
Proof.
Observe that
$\prod_{k\leq b^{\varepsilon n}}(b^{k}-1)\leq b^{b^{2\varepsilon n}}.$
If $q\leq b^{n}$ satisfies $o_{q}(b)=k\leq b^{\varepsilon n}$, then $q$ divides $b^{k}-1$. Therefore,
$b^{b^{2\varepsilon n}}\geq\prod_{\begin{subarray}{c}q\leq b^{n}\\ b\wedge q=1\\ o_{q}(b)\leq b^{\varepsilon n}\end{subarray}}q\geq N_{\varepsilon,n,b}!.$
Since for large $N$,
$N!\geq\Big(\frac{N}{e}\Big)^{N}=b^{\,N\frac{\log(N/e)}{\log b}},$
it follows, for $n$ large enough, that $N_{\varepsilon,n,b}\leq b^{2\varepsilon n}$. ∎
Write $\widetilde{Q}_{b,n}=\left\{q\in Q_{b,n}:\ q\wedge b=1\right\}.$
Combining Lemma 3.6 and Proposition 3.7, we conclude that for any $\varepsilon>0$ and all sufficiently large $n$,
$\widetilde{Q}_{b,n}\leq b^{2\varepsilon n},$
and therefore
$\lim_{n\to+\infty}\frac{\log\#\widetilde{Q}_{b,n}}{n}=0.$
Fix $q\in Q_{b,n}$ and write $q=b^{k}q^{\prime}$ where $q^{\prime}\wedge q=1.$ Fix $p\in\mathbb{N},$ $p\wedge q=1$ such that $\frac{p}{q}\in\widetilde{K}_{T}.$ Then, as $K_{T}$ is $\times b$ invariant, $\frac{b^{k}p}{q}=\frac{p}{q^{\prime}}\in\widetilde{K}_{T}$ and thus $q^{\prime}\in\widetilde{Q}_{b,n-k}.$ We conclude that, providing that $n$ is chosen large enough,
$\#Q_{b,n}\leq\sum_{0\leq k\leq n}\#\widetilde{Q}_{b,k}\leq b^{3\varepsilon n}.$
This proves Theorem 3.1.
4. Proof of Theorem 1.4
Fix $T=\{f_{1},\dots,f_{m}\}$ a self-similar IFS, let $K$ be its attractor, and let $0<c_{1}\leq\cdots\leq c_{m}<1$ be the contraction ratios of $f_{1},\dots,f_{m}$. Theorem 1.4 will be derived from the following general result.
Theorem 4.1.
Let $\eta:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be non-decreasing and assume that $T$ satisfies AWSC. Writing
$\delta_{\eta}=\liminf_{r\to 0^{+}}\frac{\log\eta(r)}{\log r},$
one has
$\dim_{H}\Big\{y\in K:\$$\big\|y-\pi\big((i_{1},\dots,i_{n}),(i_{k+1},\dots,i_{n})_{\infty}\big)\big\|_{\infty}$
$\leq\eta\Big(\Big|f_{(i_{1},\dots,i_{n})}(K)\Big|\Big)\ \text{i.o.}\Big\}=\frac{\dim_{H}K}{\max\{1,\delta_{\eta}\}},$
where i.o. means that the inequality holds for infinitely many words $(i_{1},\dots,i_{n})\in\bigcup_{p\geq 1}\{1,\dots,m\}^{p}$ and $1\leq k\leq n$.
We now explain how this applies to intrinsic Diophantine approximation.
Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be non-increasing and define
$W_{T,\psi}=\Big\{x\in K:|x-r|\leq\psi(h_{\mathrm{int}}(r))\ \text{f.i.m. }r\in\mathbb{Q}\cap K\Big\}.$
Lemma 4.2.
Setting $\eta(r)=\psi(1/r)$, the following inclusions hold:
$E_{\psi,T,\text{int}}$ $\subset\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),$
$\hskip 85.35826pt\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\Big)$
$\subset E_{\psi,T,\text{int}}\cup(\mathbb{Q}\cap K).$
Proof.
The first inclusion follows from the fact that for any $r\in\mathbb{Q}\cap K$ there exists $\underline{i}=(i_{1},\dots,i_{n})\in\{1,\dots,m\}^{n}$ such that
$h_{\mathrm{int}}(r)=\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1).$
Moreover, if
$r=\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),$
then, by monotonicity of $\eta$,
$\eta\Big(\frac{1}{h_{\mathrm{int}}(r)}\Big)\geq\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big).$
Hence, if $x$ belongs to infinitely many of the above balls, then $x\in E_{\psi,T,\text{int}}$.
Conversely, if $x$ belongs to the limsup set but only finitely many distinct rationals $r=\pi(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty})$ satisfy the corresponding inequality, then there exists $r_{x}\in\mathbb{Q}\cap K$ and infinitely many $\underline{i}$ such that
$x\in B\Big(r_{x},\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\Big),$
which implies $x=r_{x}\in\mathbb{Q}$. ∎
Corollary 4.3.
$\dim_{H}E_{\psi,T,\text{int}}=$
$\dim_{H}\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*},\,0\leq k\leq n-1}B\Big(\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\Big).$
Finally, for any $\underline{i}=(i_{1},\dots,i_{n})$,
$|f_{\underline{i}}(K)|=\frac{|K|}{\prod_{1\leq j\leq n}q_{i_{j}}}\leq\frac{|K|}{\prod_{1\leq j\leq k}q_{i_{j}}(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\leq\frac{2|K|}{\prod_{1\leq j\leq n}q_{i_{j}}}.$
Define $\widetilde{\eta},\widehat{\eta}$$\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
$\widetilde{\eta}(r)=\eta\Big(\frac{r}{|K|}\Big),\qquad\widehat{\eta}(r)=\eta\Big(\frac{2r}{|K|}\Big).$
Then $\delta_{\widetilde{\eta}}=\delta_{\widehat{\eta}}=\delta_{\eta}$ and
$\widetilde{\eta}(f_{\underline{i}}(K))\leq\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\leq\widehat{\eta}(f_{\underline{i}}(K)).$
Applying Theorem 4.1 to $\widetilde{\eta}$ and $\widehat{\eta}$ yields the desired result.
4.1. Proof of Theorem 4.1
First, note that the case $\delta_{\eta}<1$ readily follows from the case $\delta_{\eta}\geq 1$. Thus we assume throughout that $\delta_{\eta}\geq 1$.
4.1.1. Proof of the upper bound
We introduce the following notation related to the possible occurrence of exact overlaps:
(7) $\begin{cases}cl(\underline{i})=\{\underline{j}\in\Lambda^{n}:f_{\underline{j}}=f_{\underline{i}}\},\\ \mathcal{CL}_{n}=\{cl(\underline{i}):\underline{i}\in\Lambda^{n}\}.\end{cases}$
Given $\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{n}$, define
$\text{\emph{Uper}}(\underline{i})=\Big\{\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big):\underline{j}=(j_{1},\ldots,j_{n})\in cl(\underline{i}),\ 1\leq k\leq n\Big\}.$
We start with a counting lemma.
Lemma 4.4.
For every $\varepsilon>0$ there exists $N_{\varepsilon}\in\mathbb{N}$ such that for every $n\geq N_{\varepsilon}$ and any $\underline{i}\in\{1,\ldots,m\}^{n}$,
$1\leq\#\text{\emph{Uper}}(\underline{i})\leq|f_{\underline{i}}(K)|^{-\varepsilon}.$
Proof.
Set $\varepsilon_{0}=\varepsilon/5$. First observe that for any $\underline{j}\in cl(\underline{i})$ and any $1\leq k\leq n$,
$\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big)=f_{\underline{i}}\Big(\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big)\Big).$
Hence
$\text{\emph{Uper}}(\underline{i})=\#\Big\{\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big):\underline{j}\in cl(\underline{i}),\ 1\leq k\leq n\Big\}.$
Let
$n_{\underline{i}}=\Big\lfloor\frac{-\log c_{\underline{i}}}{\log 2}\Big\rfloor+1,\qquad 0\leq p\leq n_{\underline{i}}.$
Define
$\widetilde{\text{\emph{Per}}}(\underline{i},p)=\Big\{\pi((j_{k+1},\ldots,j_{n})_{\infty}):(j_{1},\ldots,j_{n})\in cl(\underline{i}),\ 1\leq k\leq n,\ 2^{-(p+1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}\Big\}.$
Note that if $f_{(\widetilde{j}_{k+1},\ldots,\widetilde{j}_{n})}=f_{(j_{k+1},\ldots,j_{n})}$, then
$\pi((\widetilde{j}_{k+1},\ldots,\widetilde{j}_{n})_{\infty})=\pi((j_{k+1},\ldots,j_{n})_{\infty}).$
Therefore,
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq\#\Big\{f_{(j_{k+1},\ldots,j_{n})}:\underline{j}\in cl(\underline{i}),\ 1\leq k\leq n,\ 2^{-(p+1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}\Big\}.$
Since for any $\underline{j}\in cl(\underline{i})$,
$2^{-n_{\underline{i}}}\leq c_{\underline{i}}=c_{\underline{j}}\leq 2^{-(n_{\underline{i}}-1)},$
we obtain
$2^{-(n_{\underline{i}}-p)}\leq c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-(n_{\underline{i}}-p-1)}.$
Moreover, if $f_{\underline{j}}(K)=f_{\underline{i}}(K)$, then
$f_{(j_{1},\ldots,j_{k})}(K)\cap f_{\underline{i}}(K)\neq\varnothing.$
By the AWSC, there exists $N_{\varepsilon_{0}}$ such that for all $n\geq N_{\varepsilon_{0}}$,
(8) $\Big\{f_{(j_{1},\ldots,j_{k})}:\underline{j}\in cl(\underline{i}),\ 1\leq k\leq n\Big\}\leq c_{\underline{i}}^{-\varepsilon_{0}}.$
Let
$\mathcal{J}_{p}=\Big\{\widetilde{j}\in cl(\underline{i}):\exists k_{\widetilde{j}}\ \text{s.t. }\ \pi((\widetilde{j}_{k_{\widetilde{j}}+1},\ldots,\widetilde{j}_{n})_{\infty})\in\widetilde{\text{\emph{Per}}}(\underline{i},p)\Big\}.$
If $f_{(j_{1},\ldots,j_{n})}(K)=f_{\underline{i}}(K)$, then
$f_{(j_{k+1},\ldots,j_{n})}(K)=f_{(j_{1},\ldots,j_{k})}^{-1}(f_{\underline{i}}(K)),$
and hence
$f_{(j_{k+1},\ldots,j_{n})}(K)\cap\bigcup_{\widetilde{j}\in\mathcal{J}_{p}}f_{(\widetilde{j}_{1},\ldots,\widetilde{j}_{k_{\widetilde{j}}})}^{-1}(f_{\underline{i}}(K))\neq\varnothing.$
For any $\widetilde{j}\in\mathcal{J}_{p}$,
$\gamma^{-1}2^{-p}\leq\Big|f_{(\widetilde{j}_{1},\ldots,\widetilde{j}_{k_{\widetilde{j}}})}^{-1}(f_{\underline{i}}(K))\Big|\leq\gamma 2^{-p},$
for some constant $\gamma\geq 1$. By (8), the above union can be covered by at most $c_{\underline{i}}^{-2\varepsilon_{0}}$ balls of radius $2^{-p}$, say $L_{1},\ldots,L_{N}$.
Since $2^{-(p-1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}$, Definition 2.4 yields, for each $1\leq\ell\leq N$,
$\Big\{f_{(j_{k+1},\ldots,j_{n})}:\pi((j_{k+1},\ldots,j_{n})_{\infty})\in\widetilde{\text{\emph{Per}}}(\underline{i},p),\ f_{(j_{k+1},\ldots,j_{n})}(K)\cap L_{\ell}\neq\varnothing\Big\}\leq c_{\underline{i}}^{-\varepsilon_{0}}.$
Consequently,
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq c_{\underline{i}}^{-3\varepsilon_{0}}.$
Summing over $0\leq p\leq n_{\underline{i}}$ gives
$\text{\emph{Uper}}(\underline{i})\leq n_{\underline{i}}c_{\underline{i}}^{-3\varepsilon_{0}}\leq c_{\underline{i}}^{-4\varepsilon_{0}}.$
Since $|f_{\underline{i}}(K)|=c_{\underline{i}}|K|$, for $n$ large enough,
$\text{\emph{Uper}}(\underline{i})\leq|f_{\underline{i}}(K)|^{-5\varepsilon_{0}}=|f_{\underline{i}}(K)|^{-\varepsilon}.$
Let $B$ be a ball and denote by $r=\operatorname{rad}(B)$ its radius. Set
$\Lambda_{B}=\Big\{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}:c_{\underline{i}}\leq r<c_{(i_{1},\ldots,i_{n-1})}\Big\}.$
Recall that $\dim_{H}K=\dim_{B}K$, where $\dim_{B}$ denotes the box dimension of $K$. Fix $0<\lambda<c_{1}$, where $c_{1}=\min_{1\leq i\leq m}c_{i}$. Let $\varepsilon>0$ and define
$s_{\varepsilon}=\frac{\dim_{H}K+2\varepsilon}{(1-\varepsilon)\delta_{\eta}}+\varepsilon.$
Choose $N_{\varepsilon}\in\mathbb{N}$ such that the following hold:
  • (1)
    Lemma 4.4 holds for all $\underline{i}\in\{1,\ldots,m\}^{n}$ with $n\geq N_{\varepsilon}$;
  • (2)
    for every $n\geq N_{\varepsilon}$ there exists a covering $\mathcal{C}_{n}$ of $K$ by balls of radius $\lambda^{n}$ satisfying
    $\mathcal{C}_{n}\leq\lambda^{-n\dim_{H}K-\varepsilon n};$
  • (3)
    for every ball $B$ with $|B|\leq\lambda^{N_{\varepsilon}}$,
    $\{f_{\underline{i}}:\underline{i}\in\Lambda_{B}\}\leq|B|^{-\varepsilon};$
  • (4)
    for any $0<r\leq\lambda^{N_{\varepsilon}}$,
    $\eta(r)\leq r^{\delta_{\eta}(1-\varepsilon)}.$
For $n\geq N_{\varepsilon}$ and $L\in\mathcal{C}_{n}$, by (1) and (3) we have
$\{f_{\underline{i}}:\underline{i}\in\Lambda_{L}\}\leq|L|^{-\varepsilon},\qquad\#\text{\emph{Uper}}(\underline{i})\leq|L|^{-\varepsilon}\quad\text{for all }\underline{i}\in\Lambda_{L},$
where we used that $\underline{i}\in\Lambda_{L}$ implies $|f_{\underline{i}}(K)|\asymp|L|$.
Moreover, since $0<\lambda<c_{1}$, for every $\underline{i}\in\bigcup_{n\geq N_{\varepsilon}}\{1,\ldots,m\}^{n}$ there exists $L\in\bigcup_{n\geq N_{\varepsilon}}\mathcal{C}_{n}$ such that $\underline{i}\in\Lambda_{L}$.
Therefore, for any $k\geq N_{\varepsilon}$,
$\limsup_{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi((i_{k+1},\ldots,i_{n})_{\infty}),\eta(|f_{\underline{i}}(K)|)\Big)$
$\subset\bigcup_{n\geq k}\bigcup_{L\in\mathcal{C}_{n}}\bigcup_{\underline{i}\in\Lambda_{L}}\bigcup_{x\in\text{\emph{Uper}}(\underline{i})}B\big(x,\eta(|f_{\underline{i}}(K)|)\big).$
Furthermore,
$\sum_{n\geq N_{\varepsilon}}\sum_{L\in\mathcal{C}_{n}}\sum_{\underline{i}\in\Lambda_{L}}\#\text{\emph{Uper}}(\underline{i})\,\eta(|L|)^{s_{\varepsilon}}$ $\leq\sum_{n\geq N_{\varepsilon}}\#\mathcal{C}_{n}\,\lambda^{-2\varepsilon n}\,\lambda^{ns_{\varepsilon}\delta_{\eta}(1-\varepsilon)}$
$\leq\sum_{n\geq N_{\varepsilon}}\lambda^{-n\dim_{H}K-2\varepsilon n+s_{\varepsilon}n\delta_{\eta}(1-\varepsilon)}$
$\leq\sum_{n\geq N_{\varepsilon}}\lambda^{n\varepsilon}<\infty.$
Consequently,
$\dim_{H}\limsup_{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi((i_{k+1},\ldots,i_{n})_{\infty}),\eta(|f_{\underline{i}}(K)|)\Big)\leq s_{\varepsilon}.$
Since this holds for every $\varepsilon>0$, we finally obtain
$\dim_{H}\limsup_{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi((i_{k+1},\ldots,i_{n})_{\infty}),\eta(|f_{\underline{i}}(K)|)\Big)\leq\frac{\dim_{H}K}{\delta_{\eta}}.$
4.1.2. Proof of the lower bound
Fix $\varepsilon>0$ and let $(r_{k})_{k\in\mathbb{N}}$ be a decreasing sequence in $\mathbb{R}_{+}^{\mathbb{N}}$ with $r_{k}\to 0$ such that, for every $k\in\mathbb{N}$,
$\eta(r_{k})\geq r_{k}^{(1+\varepsilon)\delta_{\eta}}.$
Recall that for any $k\in\mathbb{N}$,
$K=\bigcup_{\underline{i}\in\Lambda_{r_{k}}}f_{\underline{i}}(K).$
Moreover, for any $\underline{i}=(i_{1},\dots,i_{n})\in\Lambda_{r_{k}}$,
$f_{\underline{i}}(K)\subset B\Big(\pi\big((i_{1},\dots,i_{n}),(i_{n})_{\infty}\big),|f_{\underline{i}}(K)|\Big).$
This implies that
$K=\limsup_{k\in\mathbb{N},\,\underline{i}=(i_{1},\dots,i_{n})\in\Lambda_{r_{k}/c_{1}}}B\Big(\pi\big((i_{1},\dots,i_{n}),(i_{n})_{\infty}\big),\frac{r_{k}}{c_{1}}\Big).$
We recall the following result, established as [10, Proposition 6.7].
Proposition 4.5 ([10]).
There exist a self-similar IFS $T_{\varepsilon}$ and an associated self-similar measure $\mu_{\varepsilon}$ such that
$\begin{cases}\mathrm{supp}(\mu_{\varepsilon})=K,\\ \dim_{H}\mu_{\varepsilon}\geq\dim_{H}K-\varepsilon.\end{cases}$
The following mass transference principle is established in [11].
Theorem 4.6 ([11]).
Let $\mu\in\mathcal{M}(\mathbb{R})$ be a self-similar measure and let $(B_{n})_{n\in\mathbb{N}}$ be a sequence of balls centred on $\mathrm{supp}(\mu)$ such that
$\mu\Big(\limsup_{n\to+\infty}B_{n}\Big)=1.$
Then, for any $\delta\geq 1$,
$\dim_{H}\limsup_{n\to+\infty}B_{n}^{\delta}\geq\frac{\dim_{H}\mu}{\delta}.$
Since
$\mu_{\varepsilon}\Big(\limsup_{k\in\mathbb{N},\,\underline{i}=(i_{1},\dots,i_{n})\in\Lambda_{r_{k}/c_{1}}}B\Big(\pi\big((i_{1},\dots,i_{n}),(i_{n})_{\infty}\big),\frac{r_{k}}{c_{1}}\Big)\Big)=1,$
the mass transference principle yields
$\dim_{H}\limsup_{k\in\mathbb{N},\,\underline{i}=(i_{1},\dots,i_{n})\in\Lambda_{r_{k}/c_{1}}}B\Big(\pi\big((i_{1},\dots,i_{n}),(i_{n})_{\infty}\big),\Big(\frac{r_{k}}{c_{1}}\Big)^{(1+2\varepsilon)\delta_{\eta}}\Big)\geq\frac{\dim_{H}K-\varepsilon}{(1+2\varepsilon)\delta_{\eta}}.$
Note that for any $\underline{i}\in\Lambda_{r_{k}/c_{1}}$ one has $|f_{\underline{i}}(K)|\geq r_{k}$. Hence, by monotonicity of $\eta$, for all sufficiently large $k\in\mathbb{N}$,
(9) $\eta(|f_{\underline{i}}(K)|)\geq\eta(r_{k})\geq r_{k}^{\delta_{\eta}(1+\varepsilon)}\geq\Big(\frac{r_{k}}{c_{1}}\Big)^{\delta_{\eta}(1+2\varepsilon)}.$
This implies that
$\dim_{H}\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi\big((i_{1},\dots,i_{n}),(i_{k+1},\dots,i_{n})_{\infty}\big),\eta(|f_{\underline{i}}(K)|)\Big)\geq\frac{\dim_{H}K-\varepsilon}{(1+2\varepsilon)\delta_{\eta}}.$
Letting $\varepsilon\to 0$ concludes the proof of Theorem 4.1.
5. Some perspectives
5.1. Extensions of Theorem 1.1 to more general rational IFSs
We expect that similar, but more refined, tools allow one to use arguments in the same spirit as in Section 3 in order to treat IFSs of the form
$T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m},$
where $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$. More precisely, we conjecture the following.
Conjecture 5.1.
Let $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m}$, where $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, be a self-similar IFS, and let $K_{T}$ be its attractor. If $K_{T}$ has empty interior, then for any non-increasing function $\psi:\mathbb{N}\to\mathbb{R}_{+}$ one has
$\dim_{H}\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{for infinitely many }r\in\mathbb{Q}\cap K_{T}\Big\}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}},$
where $\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}$.
Regarding IFSs of the form $T=\left\{f_{i}(\cdot)=r_{i}\cdot+r^{\prime}_{i}\right\}_{1\leq i\leq m}$, where $r_{i}\in(-1,1)\cap\mathbb{Q}\setminus\{0\}$, which can always be rewritten as
$T=\left\{f_{i}(\cdot)=\frac{t_{i}\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m},$
with $q_{1},\dots,q_{m},t_{1},\dots,t_{m},p_{1},\dots,p_{m}\in\mathbb{Z}$ and $|t_{i}|<|q_{i}|$ for every $1\leq i\leq m$, we again suspect that a more refined argument in the same vein as in Section 3 may allow one to prove that, if $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, with $\delta_{\psi}$ as above,
$\dim_{H}\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{for infinitely many }r\in\mathbb{Q}\cap K_{T}\Big\}\leq\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}}.$
In this general case, the lower bound seems much more speculative. The reason is that, in general, rationals in such fractals no longer correspond to eventually periodic sequences. Thus, studying this problem requires a new description of the possible codings of rationals in this setting; see [23].
5.2. A more refined full-measure statement for homogeneous rational IFSs
Let $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q}+\frac{p_{i}}{q}\right\}$, where $q\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let $K_{T}$ denote its attractor. As mentioned in Section 2.2.2, these IFSs always satisfy the weak separation condition, and hence are Ahlfors-regular. By [22], this implies that
$\dim_{H}K_{T}=\frac{\log m}{\log q}\quad\Longleftrightarrow\quad T\text{ satisfies the open set condition}.$
In the case where $T$ satisfies the open set condition, it was proved in [24, 1] that, denoting by $\mu_{T}$ the Ahlfors-regular self-similar measure on $K_{T}$, for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$ and setting
$E_{\psi,T,\text{int}}:=\Big\{x\in K_{T}:|x-r|\leq\psi(h_{\text{int}}(r))\ \text{for infinitely many }r\in\mathbb{Q}\cap K_{T}\Big\},$
one has
$\begin{cases}\mu_{T}(E_{\psi,T,\text{int}})=1\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}=+\infty,\\[4.30554pt] \mu_{T}(E_{\psi,T,\text{int}})=0\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}<+\infty.\end{cases}$
A motivated reader can check that the argument we provide in Section 4.1.1, in this particular case (replacing the box-dimension argument by the use of the Ahlfors-regular measure), only yields that
$\sum_{n\geq 1}n^{2}m^{n}\psi(q^{n})^{\dim_{H}K_{T}}<+\infty\quad\Longrightarrow\quad\mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})=0.$
Thus, we raise the following question: is it true that
$\begin{cases}\mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})=\mathcal{H}^{\dim_{H}K_{T}}(K_{T})\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}=+\infty,\\[4.30554pt] \mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})=0\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}<+\infty?\end{cases}$
Finally, we mention that another class of self-similar IFSs known to satisfy the WSC is given by Bernoulli convolutions associated with inverses of Pisot numbers. In this particular case, although the attractor is an interval, it is conjectured (see [6]) that all self-similar measures are non-Rajchman, which implies in particular that these measures are singular with respect to Lebesgue measure. Thus, in the above statement, one might in general be able to replace the Hausdorff measure $\mathcal{H}^{\dim_{H}K_{T}}$ by a self-similar measure of full dimension.
6. Appendix
6.1. Rational numbers in translates of one-dimensional homogeneous fractals
Fix $\lambda\in(-1,1)\setminus\{0\}$ and $a_{1},\dots,a_{m}\in\mathbb{R}$, and write $T=\{f_{i}(\cdot)=\lambda\cdot+a_{i}\}_{1\leq i\leq m}$. Let $K$ be the attractor of $T$. Recall that
$K=\Big\{\sum_{k\geq 1}a_{i_{k}}\lambda^{k-1}\Big\}_{(i_{k})_{k\in\mathbb{N}}\in\{1,\dots,m\}^{\mathbb{N}}}.$
Fix $y\in\mathbb{R}$ and write $T_{y}=\{f_{i}(x)=\lambda x+a_{i}+(1-\lambda)y\}_{1\leq i\leq m}$. The attractor $K_{y}$ of $T_{y}$ is given by
$K_{y}=\Big\{\sum_{k\geq 1}(a_{i_{k}}+(1-\lambda)y)\lambda^{k-1}\Big\}_{(i_{k})_{k\in\mathbb{N}}\in\{1,\dots,m\}^{\mathbb{N}}}=y+K.$
Proposition 6.1.
[…]
Proof.
Note that $y+K\cap\mathbb{Q}\neq\emptyset$ if and only if
$y\in\bigcup_{q\in\mathbb{Q}}(q-K).$
If $\mathcal{L}(K)=0$, then
$\mathcal{L}\Big(\bigcup_{q\in\mathbb{Q}}(q-K)\Big)=0,$
which proves the second statement.
Assume now that $\mathcal{L}(K)>0$. By the Besicovitch density theorem, there exist $x\in K$ and $r_{x}>0$ such that for any $0<r\leq r_{x}$,
$\frac{\mathcal{L}(B(x,r)\cap K)}{\mathcal{L}(B(x,r))}\geq\frac{1}{2}.$
Fix $z\in\mathbb{R}$ and $r\leq r_{x}/2$. Since $\{q-x\}_{q\in\mathbb{Q}}$ is dense, there exists $q\in\mathbb{Q}$ such that $q-x\in B(z,r/2)$. Hence
$B(q-x,r/2)\cap(q-K)\subset B(z,r)\cap(q-K),$
and
$\mathcal{L}\big(B(q-x,r/2)\cap(q-K)\big)$ $=\mathcal{L}\big(B(x,r/2)\cap K\big)\geq\frac{r}{4}=\frac{1}{4}\,\mathcal{L}(B(z,r)).$
Therefore, for any $z\in\mathbb{R}$ and any $r\leq r_{x}$,
$\mathcal{L}\Big(B(z,r)\cap\bigcup_{q\in\mathbb{Q}}(q-K)\Big)\geq\frac{1}{4}\,\mathcal{L}(B(z,r)),$
which implies
$\mathcal{L}\Big(\mathbb{R}\setminus\bigcup_{q\in\mathbb{Q}}(q-K)\Big)=0.$
This concludes the proof. ∎
6.2. WSC and Ahlfors regularity
Let $T=\{f_{1},\dots,f_{m}\}$ be a self-similar IFS satisfying WSC (Definition 2.5), let $K_{T}$ be its attractor, and let $0<c_{1}\leq\cdots\leq c_{m}<1$ be the contraction ratios of $f_{1},\dots,f_{m}$.
We prove Proposition 2.3.
Fix $x\in K_{T}$ and $0<r\leq c_{1}$. Then there exists $\underline{i}\in\bigcup_{k\geq 1}\{1,\dots,m\}^{k}$ such that
$\begin{cases}f_{\underline{i}}(K)\subset B(x,r),\\ |f_{\underline{i}}(K)|\geq c_{1}r.\end{cases}$
Since $f_{\underline{i}}$ is a homothetic map with contraction ratio $c_{\underline{i}}=|f_{\underline{i}}(K)|/|K|$, recalling Theorem 2.2, we obtain
$\mathcal{H}^{\dim_{H}K_{T}}(B(x,r))\geq c_{\underline{i}}^{\dim_{H}K_{T}}\mathcal{H}^{\dim_{H}K_{T}}(K_{T})\geq\kappa_{1}r^{\dim_{H}K_{T}}.$
On the other hand, since
$K_{T}\cap B(x,r)\subset\bigcup_{\underline{i}\in\Lambda_{B(x,r)}}f_{\underline{i}}(K),$
we have
$\mathcal{H}^{\dim_{H}K_{T}}(B(x,r))$ $\leq\kappa_{2}\sup_{y\in K_{T},\rho>0}\#\Lambda_{B(y,\rho)}\,\mathcal{H}^{\dim_{H}K_{T}}(K_{T})\,r^{\dim_{H}K_{T}}$
$\leq\kappa_{3}r^{\dim_{H}K_{T}}.$